A sample of 400 racing cars showed that 80 cars cost over $700,000. What is the 99% confidence interval of the true proportion of cars costing over $700,000?
Answer by stanbon(75887)

sample of 400 racing cars showed that 80 cars cost over $700,000. What is the 99% confidence interval of the true proportion of cars costing over $700,000?
-----
p-hat = 80/400 = 1/5 = 0.2
ME = z*sqrt[p*q/n]
z = invNorm(0.995) = 2.5758
-----------------------
ME = 2.5758*sqrt[(1/5)(4/5)/400] = 0.0515
-----------
CI form:: p-hat - ME < p < p-hat + ME
-----
99% CI:: 0.2-0.0515 < p < 0.2+0.0515
------------
Cheers,
Stan H.
-----------